Find all the numbers for which the rational expression below is not defined:
x^2-16 / x^2 - 3x+12

Answer by Nate(3500)

Undefined if the denominator is zero...
x^2 - 3x + 12 = 0
Use any way to solve...
x^2 - 3x = -12
(x - 3/2)^2 = -48/4 + 9/4 = -39/4
x = 3/2 +- i*sqrt(39)/2
None exist...